OpenStudy (anonymous):

Seattle Greenery In July in Seattle, the grass grows 1/2 in. a day on a sunny day and 1/4 in. a day on a cloudy day. In Seattle in July, 75% of the days are sunny and 25% are cloudy. a) Determine the expected amount of grass growth on a typical day in July in Seattle. b) Determine the expected total grass growth in the month of July in Seattle

OpenStudy (anonymous):

75% x 0.5" + 0.25%*0.25" = 0.375 + 0.125 = 0.4375" per day July has 31 days: so 31* 0.4375 = 13.5625" We have only 2 digits of precision in the original question so we limit this to 13.56" of growth expected.